# Classification of same-gate quantum circuits and their space-time symmetries with application to the level-spacing distribution

### v1

@@ Abstract
Discussion.– We have classified all different quantum circuits in one dimension with periodic boundary conditions and translationally invariant nearest-neighbor 2-site gates. There are $N-1$ different spectral classes, being generalizations of the familiar brick-wall and the staircase configurations. Each class can be characterized by two integers $(q,r)$, such that the Floquet propagator is essentially a $q$-th power of $\tilde{F}=S^{r}f_{1}$, where for generalized S circuits one has $f_{1}=V_{1,2}$, while for generalized BW circuits $f_{1}$ is one layer of gates. We have therefore come full circle: similarly as in classical single-particle kicked models where one interchangeably applies a simple map in real space (e.g., potential $V(q)$) and a simple map in momentum space (e.g., free evolution), any quantum many-body translationally invariant Floquet system has the same basic structure. The elementary building block $\tilde{F}$ is a product of simple local transformation, like $V_{1,2}$, and of “free evolution” described by the translation operator $S$ (that is diagonal in the Fourier basis).
@@
Acknowledgements.– We acknowledge support by Grants No. J1-4385 and No. P1-0402 from the Slovenian Research Agency.
@@ Appendix A Canonical circuits
In the main text in Fig. 2 we have shown canonical circuits $F_{q,r}$ for $N=6$. Here we show in Fig. S1 all $9$ canonical $F_{q,r}$ for $N=10$, where a variety of configurations is richer. We can for instance notice that even though circuits are constructed as a $q$-layered circuit in some cases consecutive layers commute and can therefore be compressed, thus reducing the number of layers. For $N=10$ this is the case for $(q,r)=(5,2)$ and its chiral pair $(5,3)$. For $(5,2)$ the first two layers $f_{1}$ and $f_{3}$ commute and can be compressed to a single layer; likewise for the next two layers. Therefore in the compressed form $(F_{5,2})^{2}$ consists of only 5 layers instead of $10$. In this compressed form there are still $2$ idle qubits in each layer on which no gate acts. One can say that the filling fraction of gates for the circuit $(5,2)$ is $8/10$, i.e., $20\%$ of qubits is idle. The class $(5,1)$ on the other hand can not be compressed any further and has the filling fraction of only $4/10$. The only circuit with filling fraction $1$ is the standard BW with $(2,1)$.
@@ Appendix B Level-spacing distribution of a direct sum of independent RMT matrices
The general formula for the LSD of a direct sum of independent RMT matrices of arbitrary dimensions was derived in Ref. Rosenzweig and Porter (1960). The matrices in the direct sum in the theorem by Rains (Eq. (9), Ref. Rains (2003)) are approximately of equal dimensions in the large $N$ limit, which is why we use the reduced formula for the direct sum of equally dimensional matrices throughout this paper.
@@
Appendix C Proof of Theorem 1 from the main paper
@@
We will refer to $i_{l}$ as gate numbers and $l$ as time indices. When talking about the gates appearing before/after a gate $i_{l}$ in a given $F$, we will call $i_{l+1}$ the time successor and $i_{l-1}$ the time predecessor, whereas when refering to gates with neighboring numbers, we will call $i_{l}-1$ the left neighbor and $i_{l}+1$ the right neighbor of $i_{l}$. By definition, the sequences corresponding to simple circuits are permutations of the first $N$ natural numbers. For the canonical simple circuits defined in the paper in Eq. 2
@@
While the equivalence of cyclically permuted circuits was motivated by spectral equivalence, a different, perhaps more general, motivation is also possible. If we have dynamics in mind, i.e., $F^{t}$ with possibly large $t$, the definition of a starting time of our period is arbitrary, e.g., the first operator in a period could just as well have been defined as the last operator in the previous period. Therefore, it would also makes sense to define cyclically permuted circuits to be equivalent in this case.
@@ C.1 Double staircase canonical circuits $F_{p}$
We now introduce a new canonical form, different from the one in the main paper, its diagram is shown in Fig. S2. It consists of two staircase sections with opposite chirality, the second having length $p$. To show that it is indeed a canonical form, i. e. that every simple circuits is equivalent to some circuit in this form, we can prove the following Theorem.
Figure S2: Diagram of $F_{p}$ defined in Eq. (19).
Theorem S1.
@@
General step: Let’s suppose we have already shown that $F$ is equivalent to
@@
where $i^{(k)}_{l}$ are arbitrary indices relabelled in a convenient way. We can treat the sequence of gates $(1,2,\dots,k)$ as a gate, that does not commute only with $1$ and $k+1$, which means that we can repeat step 2 by trying to bring gate $k+1$ to the right of $k$ (index $k+1$). Thus, either
@@ C.2 Circuit invariant
An alternative way to determine the equivalent $F_{p}$ for a given $F$ is to calculate some quantity, which is invariant in equivalence operations and is different for all $F_{p}$. A convenient choice is the length of the second staircase $p$, which is clearly equal to the number of gates, for which their right neighbor (modulo $N$) appears in the Floquet operator before them. Let’s thus define $C(F)$ to be exactly that
@@
Let’s now show that $C(F)$ is indeed invariant under circuit equivalence operations.
Lemma S1.
@@
A transposition of time successive gates can only change $C$ if their gate number difference is ±$\pm 1$, which does not happen if the commute (Eq. (16)). Therefore, $C$ is conserved there
@@
An obvious consequence of Theorem S1 is the following lemma.
Lemma S2.
@@
According to Theorem S1, any simple circuit $F$ is equivalent to some $F_{p}$ for $p\in\{1,\dots,N-1\}$. Since $C(F)$ is conserved under equivalence transformations (Lemma S1) and $C(F_{p})=p$, $C(F)\in\{1,\dots,N-1\}$. ∎
@@ C.3 Generalized S/BW canonical circuits $F_{q,r}$
Let’s first consider the generalized S case, where $q=N$ and $\mathrm{gcd}(N,r)=1$. Because of that, it is intuitively clear, that translations by $r$ must be “ergodic” and thus generate all gates. More formally, this means that the least common multiple $\mathrm{lcm}(N,r)=rN/\mathrm{gcd}(N,r)=rN$. Since $F_{N,r}$ is generated by translating gate $1$ by $r$ sites, the lowest number of translations after which gate $1$ (modulo $N$) is generated again must be $N$ (since the smallest solution to $kr=0\pmod{N}$ is exactly $k=\mathrm{lcm}(N,r)/r$). Thus, before looping back to gate $1$ we generate $N$ gates. By an analogous argument, any gate is generated again only after $N$ translations, which means that all the generated gates are different.
Let’s now consider the generalized BW case where $q\neq N$. In this case $\mathrm{lcm}(q,N)=N$, which means that with translations by $q$ we generate $N/q$ different gates. These are exactly gates $1$ plus all the multiples of $q$ less than or equal to $N$, so gates $1+qk$, $k=0,\dots,N/q-1$. By translating them by any $r$, such that $1\leq r<q$, we generate all gates $1+r+qk$, which since $r<q$ cannot be equal to any gates generated with $r=0$. More formally, this is true because $1+r+qk\not\equiv 1+qk\pmod{q}$ necessarily implies $1+r+qk\neq 1+qk$. In order to generate all gates, $1+kr$ for $k\in\{0,1,2,\dots,q-1\}$ must now take all the possible values modulo $q$. Analogous to the generalized S case, this indeed does happen when $\mathrm{gcd}(q,r)=1$. ∎
Since Theorem S1 implies that there are $(N-1)$ non-equivalent simple circuits, a required condition for $F_{q,r}$ to be a canonical form is that there are $(N-1)$ allowed $(q,r)$.
Lemma S4.
@@
We now wish to show that the invariant $C(F_{q,r})$ is different for all allowed $(q,r)$, which is the last important statement required for the proof of Theorem 1. We do this in two steps, first for the generalized S in Lemma S5 and then for generalized BW in Lemma S6, finally combining both in Lemma S7.
Lemma S5.
@@
$C(F_{q,r})$ $=1+|\{\text{gates appearing after gate $N$}\}|$ (43)
@@
We first want to consider if we know that a certain gate is a member of $\tilde{C}(F_{N,r})$, what can we say about the membership of its time successor and predecessor. Let $F_{N,r}=(i_{1},\dots,i_{N})$. Let $i_{l}\in\tilde{C}(F_{N,r})$, which means that its right neighbor appears in $F_{N,r}$ before it, $\exists i_{k}=i_{l}+1,k<l$ (gate numbers are taken modulo $N$, but indices are not). Here $l\neq 1$, since in that case, such $k<l=1$ never exists. Let’s now consider the membership of $i_{l}$’s successor and predecessor:
@@
In other words, $C(F_{N,r})$ is coprime with $N$. According to Lemma S2, $C(F)\leq N-1$, which means that $C(F_{N,r})$ can only be a number less than $N$ and coprime with $N$
@@
Let’s now show that $k_{N}$ is different for different $F_{N,r}$ and $F_{N,\tilde{r}}$ by contradiction. Let’s suppose otherwise, this means that $k_{N}-1$ translations by $r$ and by $\tilde{r}$ must generate the same gate number modulo $N$.
@@
We have thus shown that for different allowed $r$, $C(F_{N,r})$ are different. Moreover, in Eq. (49) we have shown that the value of $C(F_{N,r})$ is a member of a set with exactly the same number of elements as allowed $r$ (see $\left|Q_{N}^{(\mathrm{S})}\right|$ in the proof of Lemma S4). This means that $C(F_{N,r})$ takes all the values from the set in Eq. (49) if we let $r$ be all the allowed $r$ values. ∎
Lemma S6.
@@
In this case, we can divide the circuit in $q$ blocks (layers) of $N/q$ time consecutive gates. The first block is generated by translating gate $1$ by $q$ (modulo $N$) until we reach gate $1$ again. The second block consists of all the gates from the first block translated by $r$. This means that the first block can be mapped to the subgroup of the cyclic group $\mathbb{Z}_{N}=\left\{0,\dots,N-1\right\}$ of order $N/q$, where group addition is taken modulo $N$. The other blocks are cosets of this subgroup (since according to Lemma S3, they must be disjoint), implying that the first $q$ gate numbers are contained in different blocks. We can therefore denote the blocks with the minimal gate number contained in it (and a tilde for clarity) $\tilde{1},\tilde{2},\dots,\tilde{q}$.
We now wish to consider a similar thing as in the proof of Lemma S5, but in the context of a block, i. e. given that a gate is a member of $\tilde{C}$, what can we say about the membership of its time successor and predecessor. We will see that the main difference is that this time, the notion time successor/predecessor can be taken inside a given block, so with indices modulo $N/q$.
Let’s suppose that a gate with index $l$ (this index is now taken in the context of a block, meaning modulo $N/q$) in block $\tilde{l}$ (modulo $q$) is a member of $\tilde{C}(F_{q,r})$, $i_{l}^{\left(\tilde{l}\right)}\in\tilde{C}(F_{q,r})$. This means that its right neighbor must appear before it, $\exists i_{k}^{\left(\tilde{k}\right)}=i_{l}^{\left(\tilde{l}\right)}+1,\tilde{k}<\tilde{l}$ (if such $k$ would exists in block $\tilde{k}=\tilde{l}$, the block would contain all $N$ gate numbers, which is not possible in the generalized BW case). Analogously to the generalized S case, $i_{l+1}^{\left(\tilde{l}\right)}=i_{l}^{\left(\tilde{l}\right)}+q$ and $i_{k+1}^{\left(\tilde{k}\right)}=i_{k}^{\left(\tilde{k}\right)}+q=i_{l}^{\left(\tilde{l}\right)}+q+1=i_{l+1}^{\left(\tilde{l}\right)}+1$ and thus $i_{l+1}^{\left(\tilde{l}\right)}\in\tilde{C}(F_{q,r})$. Crucially, indices $k$ and $l$ here are taken modulo $N/q$ because translating the last gate of the block by $q$ (modulo $N$) yields the first element of the block. Iterating this, we can conclude that if a gate from a block is contained in $\tilde{C}(F_{q,r})$, then all the gates from this block are contained in $\tilde{C}(F_{q,r})$.
We have therefore shown that by relabelling the considered circuit with $\tilde{1},\tilde{2},\dots,\tilde{q}$, the considered circuit behaves exactly the same as is in the generalized S case. Using Lemma S5, for a fixed $q$ and $N$
@@
where we took into the account that the number of gates in a block is $N/q$. Also according to Lemma S5, all the values from the set in Eq. (53) are taken for some $r$ and are different for different $r$. ∎
Lemma S7.
@@
We have calculated all the allowed values of $C(F_{q,r})$ and shown that they are different for different $r$ and a fixed $q$ in Lemma S5 and Lemma S6. The only thing left to do is to show that for different $q$, the sets (44) and (52) are disjoint.
@@
The only thing left is to check that the sets in the BW case (Eq. (52)) are disjoint for two different $q_{1},q_{2}|N$. We can do this by finding a contradiction in the converse case. Let’s suppose that we would find the same value of $C$ for different $q_{1},q_{2}$, this means that for some $p_{1},p_{2},\mathrm{gcd}(p_{1},q_{1})=\mathrm{gcd}(p_{2},q_{2})=1$ we must have
@@
Theorem 1 (restated from the paper).
@@
The fact that $F_{q,r}$ are simple circuits follows from Lemma S3 and the number of them follows from Lemma S4.
According to Lemma S7, $C(F_{q,r})$ is different for all allowed $(q,r)$ and a given $N$. Since according to Lemma S1 $C$ is conserved under equivalence transformations and, according to Eq. (29), is different for all $F_{p}$, different $F_{q,r}$ are equivalent to different $F_{p}$ (according to Theorem S1, they must be equivalent to some $F_{p}$). Moreover, since there are $(N-1)$ $F_{q,r}$ and $(N-1)$ $F_{p}$, every $F_{p}$ must be equivalent to some $F_{q,r}$, i. e. there exists a bijective equivalence mapping between $F_{q,r}$ and $F_{p}$.
Given any simple circuit $F$, according to Theorem S1, we know $F\cong F_{p}$ for some $p$. We have shown that $F_{p}\cong F_{q,r}$ for some $q,r$. Therefore, $F\cong F_{q,r}$ for some $q,r$. ∎
In contrast to the proof of Theorem S1, the proof of Theorem 1 is not constructive, i. e. it cannot be used as an algorithm to transform a given $F$ to the equivalent $F_{q,r}$. The most convenient way to do this in practice is via the invariant $C(F)$ defined in Eq. (28), which can be numerically calculated in linear time in $N$. One can then calculate $C(F)$ and $C(F_{q,r})$ for the all $(N-1)$ allowed $(q,r)$ in quadratic time. Finally, $F$ must be equivalent to the $F_{q,r}$ with the same value of $C$. As an example, the values of $C(F_{q,r})$ are included in Fig. S1.

### v2

@@ IV Level spacing statistics
If one is on the other hand interested in the eigenphases of $F_{q,r}$ one has to take into account the nontrivial modulo $2\pi$ operation. The theorem by Rains Rains (2003) tells us what is the distribution of eigenvalues of a power of a matrix from the unitary Haar measure. For $M\in\mathrm{U}_{\mathcal{N}}$, where $\mathrm{U}_{\mathcal{N}}$ denotes the Haar distribution of $\mathcal{N}\times\mathcal{N}$ unitary matrices, the eigenvalues of its power $M^{q}$ are a union of $q$ independent eigenvalue sets distributed according to $\mathrm{U}_{{\cal N}_{j}}$ of smaller matrices,
${\lambda(M^{q})\sim\bigcup_{j=0}^{q-1}\,\lambda\left(\mathrm{U}_{{\cal N}_{j}}\right),\quad{\cal N}_{j}=\left\lceil\frac{\mathcal{N}-j}{q}\right\rceil,}$ (9)
[…]
V Discussion
We have classified all different quantum circuits in one dimension with periodic boundary conditions and translationally invariant nearest-neighbor 2-site gates. There are $(N-1)$ different spectral classes, being generalizations of the familiar brick-wall and the staircase configurations. Each class can be characterized by two integers $(q,r)$, such that the Floquet propagator is essentially a $q$-th power of $\tilde{F}=S^{r}f_{1}$, where for generalized S circuits one has $f_{1}=V_{1,2}$, while for generalized BW circuits $f_{1}$ is one layer of gates. We have therefore come full circle: similarly as in classical single-particle kicked models where one interchangeably applies a simple map in real space (e.g., potential $V(q)$) and a simple map in momentum space (e.g., free evolution), any quantum many-body translationally invariant Floquet system has the same basic structure. The elementary building block $\tilde{F}$ is a product of simple local transformation, like $V_{1,2}$, and of “free evolution” described by the translation operator $S$ (that is diagonal in the Fourier basis).
@@ V Discussion
Acknowledgements.
We acknowledge support by Grants No. J1-4385 and No. P1-0402 from the Slovenian Research Agency (ARIS).
@@ Appendix A Canonical circuits
In the main text in Fig. 2 we have shown canonical circuits $F_{q,r}$ for $N=6$. Here we show in Fig. 4 all $9$ canonical $F_{q,r}$ for $N=10$, where a variety of configurations is richer. We can for instance notice that even though circuits are constructed as a $q$-layered circuit in some cases consecutive layers commute and can therefore be compressed, thus reducing the number of layers. For $N=10$ this is the case for $(q,r)=(5,2)$ and its chiral pair $(5,3)$. For $(5,2)$ the first two layers $f_{1}$ and $f_{3}$ commute and can be compressed to a single layer; likewise for the next two layers. Therefore in the compressed form $(F_{5,2})^{2}$ consists of only 5 layers instead of $10$. In this compressed form there are still $2$ idle qubits in each layer on which no gate acts. One can say that the filling fraction of gates for the circuit $(5,2)$ is $8/10$, i.e., $20\%$ of qubits is idle. The class $(5,1)$ on the other hand can not be compressed any further and has the filling fraction of only $4/10$. The only circuit with filling fraction $1$ is the standard BW with $(2,1)$.
@@ Appendix B Level-spacing distribution of a direct sum of independent RMT matrices
The general formula for the LSD of a direct sum of independent RMT matrices of arbitrary dimensions was derived in Ref. Rosenzweig and Porter (1960). The matrices in the union in the theorem by Rains Rains (2003), Eq. (9), are approximately of equal dimensions in the large $N$ limit, which is why we use the formula for the direct sum of equally dimensional matrices throughout this paper.
@@
Appendix C Proof of Theorem 1 from the main text
@@
We will refer to $i_{l}$ as gate numbers and $l$ as time indices. When talking about the gates appearing before/after a gate $i_{l}$ in a given $F$, we will call $i_{l+1}$ the time successor and $i_{l-1}$ the time predecessor, whereas when referring to gates with neighboring numbers, we will call $i_{l}-1$ the left neighbor and $i_{l}+1$ the right neighbor of $i_{l}$. By definition, the sequences corresponding to simple circuits are permutations of the first $N$ natural numbers. For the canonical simple circuits defined in the paper in Eq. 2
@@
While the equivalence of cyclically permuted circuits was motivated by spectral equivalence, a different, perhaps more general, motivation is also possible. If we have dynamics in mind, i.e., $F^{t}$ with possibly large $t$, the definition of a starting time of our period is arbitrary, e.g., the first operator in a period could just as well have been defined as the last operator in the previous period. Therefore, it would also make sense to define cyclically permuted circuits to be equivalent in this case.
@@ C.1 Double staircase canonical circuits $F_{p}$
We now introduce a new canonical form, different from the one in the main paper, its diagram is shown in Fig. 5. It consists of two staircase sections with opposite chirality, the second having length $p$. To show that it is indeed a canonical form, i. e. that every simple circuits is equivalent to some circuit in this form, we can prove the following Theorem.
Figure 5: Diagram of $F_{p}$ defined in Eq. (19).
Theorem 2.
@@
General step: Let us suppose we have already shown that $F$ is equivalent to
@@
where $i^{(k)}_{l}$ are arbitrary indices relabelled in a convenient way. We can treat the sequence of gates $(1,2,\dots,k)$ as a gate, that does not commute only with $N$ and $k+1$, which means that we can repeat step 2 by trying to bring gate $k+1$ to the right of $k$ (index $k+1$). Thus, either
@@ C.2 Circuit invariant
An alternative way to determine the equivalent $F_{p}$ for a given $F$ is to calculate some quantity, which is invariant in equivalence operations and is different for all $F_{p}$. A convenient choice is the length of the second staircase $p$, which is clearly equal to the number of gates, for which their right neighbor (modulo $N$) appears in the Floquet operator before them. Let us thus define $C(F)$ to be exactly that
@@
Let us now show that $C(F)$ is indeed invariant under circuit equivalence operations.
Lemma 1.
@@
A transposition of time successive gates can only change $C$ if their gate number difference is ±$\pm 1$, which does not happen if they commute (Eq. (16)). Therefore, $C$ is conserved there
@@
An obvious consequence of Theorem 2 is the following lemma.
Lemma 2.
@@
According to Theorem 2, any simple circuit $F$ is equivalent to some $F_{p}$ for $p\in\{1,\dots,N-1\}$. Since $C(F)$ is conserved under equivalence transformations (Lemma 1) and $C(F_{p})=p$, $C(F)\in\{1,\dots,N-1\}$. ∎
@@ C.3 Generalized S/BW canonical circuits $F_{q,r}$
Let us first consider the generalized S case, where $q=N$ and $\mathrm{gcd}(N,r)=1$. Because of that, it is intuitively clear, that translations by $r$ must be “ergodic” and thus generate all gates. More formally, this means that the least common multiple $\mathrm{lcm}(N,r)=rN/\mathrm{gcd}(N,r)=rN$. Since $F_{N,r}$ is generated by translating gate $1$ by $r$ sites, the lowest number of translations after which gate $1$ (modulo $N$) is generated again must be $N$ (since the smallest solution to $kr=0\pmod{N}$ is exactly $k=\mathrm{lcm}(N,r)/r$). Thus, before looping back to gate $1$ we generate $N$ gates. By an analogous argument, any gate is generated again only after $N$ translations, which means that all the generated gates are different.
Let us now consider the generalized BW case where $q\neq N$. In this case $\mathrm{lcm}(q,N)=N$, which means that with translations by $q$ we generate $N/q$ different gates. These are exactly gates $1$ plus all the multiples of $q$ less than or equal to $N$, so gates $1+qk$, $k=0,\dots,N/q-1$. By translating them by any $r$, such that $1\leq r<q$, we generate all gates $1+r+qk$, which since $r<q$ cannot be equal to any gates generated with $r=0$. More formally, this is true because $1+r+qk\not\equiv 1+qk\pmod{q}$ necessarily implies $1+r+qk\neq 1+qk$. In order to generate all gates, $1+kr$ for $k\in\{0,1,2,\dots,q-1\}$ must now take all the possible values modulo $q$. Analogous to the generalized S case, this indeed does happen when $\mathrm{gcd}(q,r)=1$. ∎
Since Theorem 2 implies that there are $(N-1)$ non-equivalent simple circuits, a required condition for $F_{q,r}$ to be a canonical form is that there are $(N-1)$ allowed $(q,r)$.
Lemma 4.
@@
We now wish to show that the invariant $C(F_{q,r})$ is different for all allowed $(q,r)$, which is the last important statement required for the proof of Theorem 1. We do this in two steps, first for the generalized S in Lemma 5 and then for generalized BW in Lemma 6, finally combining both in Lemma 7.
Lemma 5.
@@
$C(F_{q,r})$ $=1+|\{\text{gates appearing after gate $N$}\}|$
@@
We first want to consider if we know that a certain gate is a member of $\tilde{C}(F_{N,r})$, what can we say about the membership of its time successor and predecessor. Let $F_{N,r}=(i_{1},\dots,i_{N})$. Let $i_{l}\in\tilde{C}(F_{N,r})$, which means that its right neighbor appears in $F_{N,r}$ before it, $\exists i_{k}=i_{l}+1,k<l$ (gate numbers are taken modulo $N$, but indices are not). Here $l\neq 1$, since in that case, such $k<l=1$ never exists. Let us now consider the membership of $i_{l}$’s successor and predecessor:
@@
In other words, $C(F_{N,r})$ is coprime with $N$. According to Lemma 2, $C(F)\leq N-1$, which means that $C(F_{N,r})$ can only be a number less than $N$ and coprime with $N$
@@
Let us now show that $k_{N}$ is different for different $F_{N,r}$ and $F_{N,\tilde{r}}$ by contradiction. Let us suppose otherwise, this means that $k_{N}-1$ translations by $r$ and by $\tilde{r}$ must generate the same gate number modulo $N$.
@@
We have thus shown that for different allowed $r$, $C(F_{N,r})$ are different. Moreover, in Eq. (47) we have shown that the value of $C(F_{N,r})$ is a member of a set with exactly the same number of elements as allowed $r$ (see $\left|Q_{N}^{(\mathrm{S})}\right|$ in the proof of Lemma 4). This means that $C(F_{N,r})$ takes all the values from the set in Eq. (47) if we let $r$ be all the allowed $r$ values. ∎
Lemma 6.
@@
In this case, we can divide the circuit in $q$ blocks (layers) of $N/q$ time consecutive gates. The first block is generated by translating gate $1$ by $q$ (modulo $N$) until we reach gate $1$ again. The second block consists of all the gates from the first block translated by $r$. This means that the first block can be mapped to the subgroup of the cyclic group $\mathbb{Z}_{N}=\left\{0,\dots,N-1\right\}$ of order $N/q$, where group addition is taken modulo $N$. The other blocks are cosets of this subgroup (since according to Lemma 3, they must be disjoint), implying that the first $q$ gate numbers are contained in different blocks. We can therefore denote the blocks with the minimal gate number contained in it (and a tilde for clarity) $\tilde{1},\tilde{2},\dots,\tilde{q}$.
We now wish to consider a similar thing as in the proof of Lemma 5, but in the context of a block, i. e. given that a gate is a member of $\tilde{C}$, what can we say about the membership of its time successor and predecessor. We will see that the main difference is that this time, the notion time successor/predecessor can be taken inside a given block, so with indices modulo $N/q$.
Let us suppose that a gate with index $l$ (this index is now taken in the context of a block, meaning modulo $N/q$) in block $\tilde{l}$ (modulo $q$) is a member of $\tilde{C}(F_{q,r})$, $i_{l}^{\left(\tilde{l}\right)}\in\tilde{C}(F_{q,r})$. This means that its right neighbor must appear before it, $\exists i_{k}^{\left(\tilde{k}\right)}=i_{l}^{\left(\tilde{l}\right)}+1,\tilde{k}<\tilde{l}$ (if such $k$ would exists in block $\tilde{k}=\tilde{l}$, the block would contain all $N$ gate numbers, which is not possible in the generalized BW case). Analogously to the generalized S case, $i_{l+1}^{\left(\tilde{l}\right)}=i_{l}^{\left(\tilde{l}\right)}+q$ and $i_{k+1}^{\left(\tilde{k}\right)}=i_{k}^{\left(\tilde{k}\right)}+q=i_{l}^{\left(\tilde{l}\right)}+q+1=i_{l+1}^{\left(\tilde{l}\right)}+1$ and thus $i_{l+1}^{\left(\tilde{l}\right)}\in\tilde{C}(F_{q,r})$. Crucially, indices $k$ and $l$ here are taken modulo $N/q$ because translating the last gate of the block by $q$ (modulo $N$) yields the first element of the block. Iterating this, we can conclude that if a gate from a block is contained in $\tilde{C}(F_{q,r})$, then all the gates from this block are contained in $\tilde{C}(F_{q,r})$.
We have therefore shown that by relabelling the considered circuit with $\tilde{1},\tilde{2},\dots,\tilde{q}$, the considered circuit behaves exactly the same as is in the generalized S case. Using Lemma 5, for a fixed $q$ and $N$
@@
where we took into the account that the number of gates in a block is $N/q$. Also according to Lemma 5, all the values from the set in Eq. (51) are taken for some $r$ and are different for different $r$. ∎
Lemma 7.
@@
We have calculated all the allowed values of $C(F_{q,r})$ and shown that they are different for different $r$ and a fixed $q$ in Lemma 5 and Lemma 6. The only thing left to do is to show that for different $q$, the sets (43) and (50) are disjoint.
@@
The only thing left is to check that the sets in the BW case (Eq. (50)) are disjoint for two different $q_{1},q_{2}|N$. We can do this by finding a contradiction in the converse case. Let us suppose that we would find the same value of $C$ for different $q_{1},q_{2}$, this means that for some $p_{1},p_{2},\mathrm{gcd}(p_{1},q_{1})=\mathrm{gcd}(p_{2},q_{2})=1$ we must have
@@
Theorem 1 (restated from the main text).
@@
The fact that $F_{q,r}$ are simple circuits follows from Lemma 3 and the number of them follows from Lemma 4.
According to Lemma 7, $C(F_{q,r})$ is different for all allowed $(q,r)$ and a given $N$. Since according to Lemma 1 $C$ is conserved under equivalence transformations and, according to Eq. (29), is different for all $F_{p}$, different $F_{q,r}$ are equivalent to different $F_{p}$ (according to Theorem 2, they must be equivalent to some $F_{p}$). Moreover, since there are $(N-1)$ $F_{q,r}$ and $(N-1)$ $F_{p}$, every $F_{p}$ must be equivalent to some $F_{q,r}$, i. e. there exists a bijective equivalence mapping between $F_{q,r}$ and $F_{p}$.
Given any simple circuit $F$, according to Theorem 2, we know $F\cong F_{p}$ for some $p$. We have shown that $F_{p}\cong F_{q,r}$ for some $q,r$. Therefore, $F\cong F_{q,r}$ for some $q,r$. ∎
In contrast to the proof of Theorem 2, the proof of Theorem 1 is not constructive, i. e. it cannot be used as an algorithm to transform a given $F$ to the equivalent $F_{q,r}$. The most convenient way to do this in practice is via the invariant $C(F)$ defined in Eq. (28), which can be numerically calculated in linear time in $N$. One can then calculate $C(F)$ and $C(F_{q,r})$ for the all $(N-1)$ allowed $(q,r)$ in quadratic time. Finally, $F$ must be equivalent to the $F_{q,r}$ with the same value of $C$. As an example, the values of $C(F_{q,r})$ are included in Fig. 4.